# The $\alpha$–regression for compositional data: a unified framework for standard, spatially-lagged, spatial autoregressive and geographically-weighted regression models

### v2

@@ Abstract
Compositional data—vectors of non-negative components summing to unity—frequently arise in scientific applications where covariates influence the relative proportions of components, yet traditional regression approaches ace challenges regarding the unit-sum constraint and zero values. This paper revisits the $\alpha$–regression framework, which uses a flexible power transformation parameterized by $\alpha$ to interpolate between raw data analysis and log-ratio methods, naturally handling zeros without imputation while allowing data-driven transformation selection. We formulate $\alpha$–regression as a non-linear least squares problem, provide efficient estimation via the Levenberg-Marquardt algorithm, and derive marginal effects for interpretation. The framework is extended to spatial settings through two models: the $\alpha$–spatially lagged X regression model, which incorporates spatial spillover effects via spatially lagged covariates with decomposition into direct and indirect effects, the $\alpha$–spatially autoregressive regression model and the geographically weighted $\alpha$–regression, which allows coefficients to vary spatially for capturing local relationships. Applications to two real data sets illustrate the performance of the models and showcase that spatial extensions capture the spatial dependence and improve the predictive performance.
@@ 1 Introduction
Compositional data are vectors of non-negative components summing to a constant, typically equal 1, for simplicity. Their sample space is the standard simplex
@@
where $D$ denotes the number of variables (components), and d=D1d=D-1.
Examples of compositional data may be found in many different fields of study and the extensive scientific literature that has been published on the proper analysis of this type of data is indicative of its prevalence in real-life applications [^1] For a substantial number of specific examples of applications involving compositional data see (tsagris2020)..
It is unsurprising, considering the ubiquity of compositional data, that many applications of compositional data analysis incorporate covariates. Examples include glacial compositional data, household consumption expenditures, concentrations of chemical elements in soil samples, morphometric fish measurements, as well as data on elections, pollution, and energy, all of which are associated with some covariates. Beyond these cases, the literature provides numerous further applications of compositional regression. For example, oceanography research involving Foraminiferal compositions at various sea depths was analyzed in ait2003. In hydrochemistry, regression methods were used by otero2005 to distinguish anthropogenic from geological sources of river pollution in Spain. Economic studies such as morais2018 connected market shares to covariates, while political science research linked candidate vote percentages to relevant predictors (katz1999). In bioinformatics, compositional approaches have also been applied to microbiome data analysis (xia2013, chen2016, shi2016).
The practical demand for robust regression models tailored to compositional data has led to numerous methodological advances, especially in recent years. The first such model was introduced by ait2003—commonly known as Aitchison’s model—based on log-ratio transformations, yielding the log-ratio approach (LRA). ilr2003 advanced Aitchison’s model by applying an isometric log-ratio transformation. The stay-in-the-simplex approach on the other hand employs distributions and models defined on the simplex. Dirichlet regression for instance has been employed in compositional contexts gueorguieva2008, hijazi2009, melo2009. Moreover, iyengar2002 examined the generalized Liouville distribution family, which allows negative or mixed correlations and extends beyond Dirichlet distributions to include non-positive correlation structures. A less common approach is to ignore the compositional constraint and treat the data as though they were Euclidean, an approach termed raw data analysis (RDA) (baxter2001, baxter2005). A fourth approach is to employ a general family of transformations, namely the $\alpha$–transformation (tsagris2011) that interpolates between the and the RDA and the LRA, offers a higher flexibility and treats zero values naturally.
A limitation of the regression models discussed above is their inability to directly accommodate zero values. As a result, several models have been developed more recently to resolve this limitation. For instance, scealy2011 mapped compositional data onto the unit hyper-sphere and proposed the Kent regression, which naturally accounts for zeros. From a Bayesian perspective, spatial compositional data containing zeros were modeled in leininger2013. In the context of economics, mullahy2015 estimated regression models for share data where the proportions could assume zero values with non-negligible probability. Further econometric approaches suitable for handling zeros are reviewed in murteira2016. In addition, tsagris2015a introduced a regression framework based on minimizing the Jensen–Shannon divergence. tsagris2018 extended Dirichlet regression to allow zeros, resulting in what is termed zero-adjusted Dirichlet regression. More recently, alenazi2022 studied and examined the properties of the $\phi$-divergence regression models, which are suitable for compositional data with zeros.
When it comes spatial autocorrelation models, a simple version is the spatial distributed lag model with spatial lags on covariates, commonly known as the spatially lagged X (SLX) model. Unlike the general spatial Durbin or spatial autoregressive models, the SLX model incorporates spatial dependence only through the covariates, excluding the spatial lag of the dependent variable (lesage2009introduction, elhorst2014spatial).
Another model is the spatial autoregressive (SAR) model that similarly to the autoregressive model in time series it assumed that observations are influenced by those in nearby locations. In the SAR model, the dependent variable at each location is expressed as a function of both explanatory variables and a weighted average of the dependent variable values from neighboring areas. nguyen2021 proposed a SAR model that employs the ilr transformation.
A local form of linear regression, used to model spatially varying relationships, is the geographically weighted regression (GWR) is. Unlike traditional regression which assumes stationarity in the relationship between dependent and independent variables, GWR allows model parameters to vary over space. The integration of GWR with compositional data analysis is relatively recent. One key challenge is reconciling the spatial non-stationarity modeled by GWR with the constraints inherent in compositional data. Several approaches have been proposed. leininger2013 combined GWR with hierarchical Bayesian frameworks for compositional data with zero values, allowing for spatial priors that account for local variation. yoshida2021 applied the isometric log-ratio (ilr) transformation before applying GWR. This preserves the relative information between parts while enabling spatially varying coefficient estimation. Finally, clarotto2022 introduced a new power transformation, similar in spirit to the $\alpha$–transformation, for geostatistical modeling of compositional data.
@@ 2.1 The $\alpha$–transformation
tsagris2011 introduced the $\alpha$–transformation, a power-based mapping designed for compositional data, $\mathbf{y}=(y_{1},y_{2},\ldots,y_{D})$. For a given parameter $\alpha\in[-1,1]$, the transformation is defined in two steps. Each component is raised to the power $\alpha$ and renormalized to remain in the simplex
@@ 2.2 The $\alpha$–regression
tsagris2015b used the log-likelihood of the multivariate normal distribution, but in this paper the regression is formulated as a non-linear least squares problem, where the minimizing function is the sum of squares of the errors (SSE)
@@
$\frac{\mu_{i}^{\alpha}}{\sum_{j=1}^{D}\mu_{j}^{\alpha}}=\frac{\left(\frac{e^{{\bf x}^{\top}\boldsymbol{\beta}_{i}}}{1+\sum_{j=1}^{D}e^{{\bf x}^{\top}\boldsymbol{\beta}_{j}}}\right)^{\alpha}}{\frac{1+\sum_{k=1}^{D}\left(e^{{\bf x}^{\top}\boldsymbol{\beta}_{k}}\right)^{\alpha}}{\left(1+\sum_{j=1}^{D}e^{{\bf x}^{\top}\boldsymbol{\beta}_{j}}\right)^{\alpha}}}=\frac{\left(e^{{\bf x}^{\top}\boldsymbol{\beta}_{i}}\right)^{\alpha}}{1+\sum_{j=1}^{D}\left(e^{{\bf x}^{\top}\boldsymbol{\beta}_{i}}\right)^{\alpha}}.$
For a given value of $\alpha$, the matrix of the regression coefficients $\mathbf{B}=\left(\mathbf{\beta}_{1},\ldots,\mathbf{\beta}_{d}\right)$ is estimated using a modification of the Levenberg-Marquardt algorithm [^2] This algorithm interpolates between the Gauss–Newton algorithm and the method of gradient descent.. The R package minpack.lm (minpacklm2023) is employed to this end [^3] The relevant gradient vector, and the Hessian matrix are provided in the Appendix. The Newton-Raphson algorithm was also tested but it is slower..
@@ 2.2.1 Limiting case of $\alpha\rightarrow 0$
tsagris2016 presented the proof that as $\alpha\rightarrow 0$, the $\alpha$–transformation (3) converges to the ilr transformation (5). Following similar calculations one can show that
@@ 2.2.2 Choosing $\alpha$
In the regression setting the optimal value of $\alpha$ is data-driven, and there are two ways to estimate its value. The first is minimize the Kullback-Leibler divergence (KLD) between the observed and fitted compositions $\text{KLD}\left(\mathbf{y},\mathbf{\mu}\right)=\sum_{i=1}^{n}\sum_{j=1}^{D}\mathbf{y}_{ij}\log{y_{i}j/\mu_{ij}}$. This results in a double minimization problem. For a given value of $\alpha$ one must minimize the SSE (7) in order to obtain the regression coefficients and then minimize the KLD with respect to $\alpha$ to obtain the optimal value of $\alpha$. This way and with the choice of the KLD, the value of $\alpha$ is independent of the SSE, since the SSE is not comparable across the different values of $\alpha$.
The second option is to examine $\alpha$ as a hyper-parameter whose value is chosen by minimizing the KLD via CV, e.g. 10-fold CV. (tsagris2015b).
@@ 2.2.4 Standard error of the MEs
First, we need the derivatives of $\mu$ with respect to $\mathbf{\beta}$:
@@ 2.2.5 Advantages and limitations of the $\alpha$–regression
The advantages of the $\alpha$–regression are: a) ability to handle zeros naturally without imputation. b) Flexible, as $\alpha$ provides a continuum from power transforms to log-ratio methods. c) Often yields better predictive performance than classical methods. d) This method balances the strengths of power transformations and log-ratio methods, providing a flexible and effective tool for predictive modeling on the simplex. Disadvantages on the other hand are a) the interpretability of regression coefficients is reduced compared to log-ratio approaches. b) The focus is mainly on prediction rather than inference; theoretical properties of estimators have not been rigorously developed.
2.3 Extension of the $\alpha$–regression and related models

### v3

@@ Abstract
Compositional data-vectors of non-negative components summing to unity-frequently arise in scientific applications where covariates influence the relative proportions of components, yet traditional regression approaches ace challenges regarding the unit-sum constraint and zero values. This paper revisits the $\alpha$–regression framework, which uses a flexible power transformation parameterized by $\alpha$ to interpolate between raw data analysis and log-ratio methods, naturally handling zeros without imputation while allowing data-driven transformation selection. We formulate $\alpha$–regression as a non-linear least squares problem, provide efficient estimation via the Levenberg-Marquardt algorithm, and derive marginal effects for interpretation. The framework is extended to spatial settings through two models: the $\alpha$–spatially lagged X regression model, which incorporates spatial spillover effects via spatially lagged covariates with decomposition into direct and indirect effects, the $\alpha$–spatially autoregressive regression model and the geographically weighted $\alpha$–regression, which allows coefficients to vary spatially for capturing local relationships. Applications to two real data sets illustrate the performance of the models and showcase that spatial extensions capture the spatial dependence and improve the predictive performance.
@@ 1 Introduction
Compositional data are characterized as vectors of non-negative components constrained to sum to a constant, conventionally normalized to unity. The sample space of such data is defined by the standard simplex
@@
where $D$ denotes the dimensionality of the compositional vector, and d=D1d=D-1 represents the degrees of freedom.
Compositional data structures arise across diverse scientific domains, as evidenced by the substantial body of methodological literature devoted to their rigorous statistical analysis [^1] For an extensive compilation of domain-specific applications involving compositional data, see (Tsagris and Stewart, 2020)..
The methodological imperative to develop models specifically calibrated for compositional data has catalyzed considerable innovation, particularly in the contemporary statistical literature. The foundational framework was established by Aitchison (2003)—subsequently designated as Aitchison’s model—predicated upon log-ratio transformations, thereby inaugurating the log-ratio approach (LRA). This methodology was subsequently refined by Egozcue et al. (2003), who implemented an isometric log-ratio transformation to preserve geometric properties. In contrast, the stay-in-the-simplex approach employs probability distributions and regression structures intrinsically defined on the simplex manifold. Notably, Dirichlet regression has been extensively utilized within compositional frameworks Gueorguieva et al. (2008), Hijazi and Jernigan (2009), Melo et al. (2009). Furthermore, Iyengar and Dey (2002) investigated the generalized Liouville distribution family, which accommodates negative or heterogeneous correlation structures, thereby extending beyond the restrictive positive correlation constraint of Dirichlet distributions. A less theoretically justified, yet occasionally employed strategy involves disregarding the unit-sum constraint and treating compositional data within a Euclidean framework—an approach designated as raw data analysis (RDA) (Baxter, 2001, Baxter et al., 2005). A fourth methodological paradigm employs the $\alpha$–transformation family (Tsagris et al., 2011), which interpolates continuously between the RDA and LRA, thereby affording enhanced model flexibility while accommodating zero components naturally.
Given the ubiquity of compositional data across scientific disciplines, regression frameworks incorporating covariates have become methodologically essential. Applications encompass glacial sediment compositions, household expenditure allocations, geochemical soil analyses, morphometric measurements, electoral outcomes, pollution indices, and energy consumption patterns, among numerous other domains. The literature documents extensive applications of compositional regression methodology. For instance, Aitchison (2003) analyzed foraminiferal compositions across bathymetric gradients in oceanographic research. In hydrochemistry, Otero et al. (2005) employed regression techniques to discriminate anthropogenic from lithogenic sources of riverine contamination in Spain. Economic research, exemplified by Morais et al. (2018), has related market share dynamics to explanatory variables, while political scientists have modeled candidate vote proportions as functions of demographic and socioeconomic predictors (Katz and King, 1999). Compositional methodologies have similarly been deployed in bioinformatic analyses of microbiome community structures (Xia et al., 2013, Chen and Li, 2016, Shi et al., 2016).
A fundamental limitation of the aforementioned regression models concerns their inability to accommodate zero-valued components directly. Consequently, methodological developments have addressed this constraint through various approaches. Scealy and Welsh (2011) proposed a transformation mapping compositional data to the unit hypersphere, introducing Kent regression which naturally accommodates structural zeros. From a Bayesian hierarchical perspective, Leininger et al. (2013) developed models for spatially indexed compositional data containing zeros. Within econometric contexts, Mullahy (2015) formulated regression frameworks for fractional response data exhibiting zero inflation. Additional econometric strategies suitable for zero-augmented compositional data are systematically reviewed in Murteira and Ramalho (2016). Moreover, Tsagris (2015a) introduced a regression framework predicated upon minimization of the Jensen–Shannon divergence. Tsagris and Stewart (2018) extended the Dirichlet regression paradigm to accommodate zero components, yielding zero-adjusted Dirichlet regression. More recently, Alenazi (2022) investigated the properties of $\phi$-divergence regression models applicable to zero-inflated compositional data.
Concerning spatial autocorrelation structures, the spatially lagged X (SLX) model represents a parsimonious specification incorporating spatial dependence exclusively through exogenous covariates, thereby excluding spatial lags of the dependent variable (LeSage and Pace, 2009, Elhorst, 2014). The spatial autoregressive (SAR) model, analogous to temporal autoregressive processes, posits that observations are influenced by proximate spatial neighbors. Specifically, the SAR model expresses the dependent variable as a function of both explanatory covariates and a spatially weighted average of neighboring dependent variable realizations. Nguyen et al. (2021) developed a SAR specification employing the ilr transformation for compositional responses. Geographically weighted regression (GWR) constitutes a local regression methodology designed to capture spatially heterogeneous relationships. In contrast to conventional regression, which assumes parameter stationarity, GWR permits spatial nonstationarity through location-specific coefficient estimation.
The integration of the spatial regression framework within the compositional data analysis represents a relatively narrow research area. Leininger et al. (2013) synthesized GWR with hierarchical Bayesian models for zero-inflated compositional data, incorporating spatial priors to accommodate local variation. Yoshida et al. (2021) implemented the isometric log-ratio transformation prior to GWR estimation, thereby preserving inter-component relationships while facilitating spatially varying parameter estimation. Recently, Clarotto et al. (2022) introduced a novel power transformation, conceptually analogous to the $\alpha$–transformation, specifically calibrated for geostatistical modeling of compositional data.
This paper adopts a pragmatic methodological stance, particularly tailored to regression with compositional data. The principal contribution of this paper is a unified framework for regression modelling of compositional data. The $\alpha$–regression (Tsagris, 2015b), was proposed as a generalization of Aitchison’s log-ratio regression (Aitchison, 2003) that naturally accommodates zero components, while offering flexibility via the $\alpha$–transformation. Based on this, extensions to accomodate spatial correlations are developed.
@@ 2.1 The $\alpha$–transformation
Tsagris et al. (2011) introduced the $\alpha$–transformation, a power-based mapping designed for compositional data, $\mathbf{y}=(y_{1},y_{2},\ldots,y_{D})$. For a given parameter $\alpha\in[-1,1]$, the transformation is defined in two steps. Each component is raised to the power $\alpha$ and renormalized to remain in the simplex
@@ 2.2 The $\alpha$–regression
Tsagris (2015b) used the log-likelihood of the multivariate normal distribution, but in this paper the regression is formulated as a non-linear least squares problem, where the minimizing function is the sum of squares of the errors (SSE)
@@
$\frac{\mu_{i,\alpha}}{\sum_{j=1}^{D}\mu_{j,\alpha}}=\frac{\left(\frac{e^{{\bf x}^{\top}\boldsymbol{\beta}_{i}}}{1+\sum_{j=1}^{D}e^{{\bf x}^{\top}\boldsymbol{\beta}_{j}}}\right)^{\alpha}}{\frac{1+\sum_{k=1}^{D}\left(e^{{\bf x}^{\top}\boldsymbol{\beta}_{k}}\right)^{\alpha}}{\left(1+\sum_{j=1}^{D}e^{{\bf x}^{\top}\boldsymbol{\beta}_{j}}\right)^{\alpha}}}=\frac{\left(e^{{\bf x}^{\top}\boldsymbol{\beta}_{i}}\right)^{\alpha}}{1+\sum_{j=1}^{D}\left(e^{{\bf x}^{\top}\boldsymbol{\beta}_{i}}\right)^{\alpha}}.$
For a given value of $\alpha$, the matrix of the regression coefficients $\mathbf{B}=\left(\mathbf{\beta}_{1},\ldots,\mathbf{\beta}_{d}\right)$ is estimated using a modification of the Levenberg-Marquardt algorithm [^2] This algorithm interpolates between the Gauss–Newton algorithm and the method of gradient descent.. The R package minpack.lm (Elzhov et al., 2023) is employed to this end [^3] The relevant gradient vector, and the Hessian matrix are provided in the Appendix. The Newton-Raphson algorithm was also tested but it is slower..
@@ 2.2.1 Limiting case of $\alpha\rightarrow 0$
Tsagris et al. (2016) presented the proof that as $\alpha\rightarrow 0$, the $\alpha$–transformation (3) converges to the ilr transformation (5). Following similar calculations one can show that
@@ 2.2.2 Choosing $\alpha$
In the regression setting the optimal value of $\alpha$ is data-driven, and there are two ways to estimate its value. The first is to minimize the Kullback-Leibler divergence (KLD) between the observed and fitted compositions $\text{KLD}\left(\mathbf{y},\mathbf{\mu}\right)=\sum_{i=1}^{n}\sum_{j=1}^{D}\mathbf{y}_{ij}\log{y_{i}j/\mu_{ij}}$. This results in a double minimization problem. For a given value of $\alpha$ one must minimize the SSE (7) in order to obtain the regression coefficients and then minimize the KLD with respect to $\alpha$ to obtain the optimal value of $\alpha$. With the choice of the KLD, the value of $\alpha$ is independent of the SSE, since the SSE is not comparable across the different values of $\alpha$.
The second option is to examine $\alpha$ as a hyper-parameter whose value is chosen by minimizing the KLD via CV, e.g. 10-fold CV. (Tsagris, 2015b).
@@ 2.2.4 Standard error of the MEs
The derivatives of $\mu$ with respect to $\mathbf{\beta}$ are:
@@ 2.2.5 Advantages and limitations of the $\alpha$–regression
The advantages of the $\alpha$–regression are: a) ability to handle zeros naturally without imputation. b) Flexibility as $\alpha$ provides a continuum from power transforms to log-ratio methods. c) A predictive performance than is often higher compared to classical methods. d) A balance of the strengths of power transformations and log-ratio methods, providing a flexible and effective tool for predictive modeling on the simplex. A disadvantage though is the reduced interpretability of regression coefficients compared to log-ratio approaches.
2.3 Consistency and asymptotic normality
To establish the large-sample properties of the $\alpha$-regression estimator $\hat{\theta}_{n}$, we impose the following assumptions or regularity conditions. The four structural and consistency assumptions delineated below are necesary for the consistency.
Assumption 2.1 (Data Generating Process).
The data $\{(y_{i},x_{i})\}_{i=1}^{n}$ are independent and identically distributed (i.i.d.), and the conditional mean is correctly specified
$\mathbb{E}[y_{i,\alpha}\mid x_{i}]=\mu_{i,\alpha}(\theta_{0})\,.$
Assumption 2.2 (Parameter Space).
The parameter space $\Theta\subset\mathbb{R}^{d(p+1)}$ is compact, and the true parameter $\theta_{0}\in\text{int}(\Theta)$.
Assumption 2.3 (Identification).
The population objective function
$Q(\theta)=\mathbb{E}[\|y_{i,\alpha}-\mu_{i,\alpha}(\theta)\|^{2}]$
has a unique minimum at $\theta_{0}$. That is, for all $\theta\neq\theta_{0}$, $Q(\theta)>Q(\theta_{0})$.
Assumption 2.4 (Continuity and Moment Conditions).
The function $\mu_{i,\alpha}(\theta)$ is continuous in $\theta$ for all $\theta\in\Theta$. Furthermore,
$\mathbb{E}[\|y_{i,\alpha}\|^{2}]<\infty\,\,\,\,\,\,\text{and}\,\,\,\,\,\,\mathbb{E}[\sup_{\theta\in\Theta}\|\mu_{i,\alpha}(\theta)\|^{2}]<\infty\,.$
Theorem 2.5 (Consistency).
Under Assumptions 2.12.4,
$\hat{\theta}_{n}\xrightarrow{p}\theta_{0}\,\,\,\,\,\,\text{as}\,\,\,\,\,\,n\to\infty\,.$
Remark 2.6 (Necessity of Correct Specification in Consistency).
The assumption $\mathbb{E}[y_{i,\alpha}\mid x_{i}]=\mu_{i,\alpha}(\theta_{0})$ (Assumption 2.1) ensures that the population objective function $Q(\theta)=\mathbb{E}[\left\|y_{i,\alpha}-\mu_{i,\alpha}(\theta)\right\|^{2}]$ is globally minimized at the true parameter $\theta_{0}$. Under correct specification, the first-order condition at $\theta_{0}$ becomes $\mathbb{E}[g_{i}(\theta_{0})^{\top}(y_{i,\alpha}-\mu_{i,\alpha}(\theta_{0}))]=0$. If the mean is misspecified, the estimator $\hat{\theta}_{n}$ will converge to a value $\theta^{*}$ that minimizes the approximation error, leading to asymptotic bias.
Remark 2.7 (Necessity of Bound on $y_{i,\alpha}$).
Although the raw composition $y_{i}$ is bounded within the simplex $\mathcal{S}^{D-1}$, the transformed response $y_{i,\alpha}$ is not necessarily bounded. Specifically, for $\alpha<0$, $y_{i,\alpha}$ involves terms of the form $y_{ij}^{-|\alpha|}$, which diverge as any component approaches zero. The assumption $\mathbb{E}[\|y_{i,\alpha}\|^{2}]<\infty$ is therefore essential to ensure that the objective function has a finite population expectation and that the variance of the residuals remains manageable.
Remark 2.8 (Dominance and Uniform Convergence).
The assumption $\mathbb{E}[\sup_{\theta\in\Theta}\|\mu_{i,\alpha}(\theta)\|^{2}]<\infty$ acts as a dominance condition. For the estimator $\hat{\theta}_{n}$ to be consistent, we require the sample objective $Q_{n}(\theta)$ to converge to the population objective $Q(\theta)$ uniformly over $\Theta$. This moment condition ensures that the random objective function is bounded by an integrable function, satisfying the requirements for the USLLN.
Remark 2.9 (Identification and Compactness).
Consistency requires the identification of $\theta_{0}$ as a well-separated minimum. By assuming $\Theta$ is compact and $\mu_{i,\alpha}(\theta)$ is continuous, the existence of a minimum is guaranteed. The moment conditions further ensure that the surface of the objective function is sufficiently ”smooth” in expectation, preventing the estimator from converging to boundary points or spurious local minima as $n\to\infty$.
The next three assumptions are necessary for asymptotic normality.
Assumption 2.10 (Smoothness).
The fitted values $\mu_{i,\alpha}(\theta)$ are twice continuously differentiable in $\theta$ on $\text{int}(\Theta)$. Let $g_{i}(\theta)=-\partial\mu_{i,\alpha}(\theta)/\partial\theta$ be the Jacobian.
Assumption 2.11 (Strengthened Moment Conditions).
The following hold: $\mathbb{E}[\sup_{\theta\in\Theta}\|g_{i}(\theta)\|^{2}]<\infty$ and
$\mathbb{E}[\sup_{\theta\in\Theta}\|\nabla^{2}r_{i}(\theta)\|]<\infty$.
Assumption 2.12 (Non-singularity).
The Gram matrix $G(\theta_{0})=\mathbb{E}[g_{i}(\theta_{0})^{\top}g_{i}(\theta_{0})]$ is positive definite.
Based on the regularity conditions established previously[cite: 11, 13, 17, 18], we now prove that the estimator $\hat{\theta}_{n}$ is asymptotically normally distributed.
Theorem 2.13 (Asymptotic Normality).
Under Assumptions 2.1-2.12, as $n\to\infty$:
$\sqrt{n}(\hat{\theta}_{n}-\theta_{0})\xrightarrow{d}N(0,G(\theta_{0})^{-1}\Omega(\theta_{0})G(\theta_{0})^{-1})$ (14)
where $G(\theta_{0})=\mathbb{E}[g_{i}(\theta_{0})^{\top}g_{i}(\theta_{0})]$ and $\Omega(\theta_{0})=\mathbb{E}[g_{i}(\theta_{0})^{\top}r_{i}(\theta_{0})r_{i}(\theta_{0})^{\top}g_{i}(\theta_{0})]$.
Remark 2.14 (Necessity of Correct Specification in Asymptotic Normality).
The assumption $\mathbb{E}[y_{i,\alpha}\mid x_{i}]=\mu_{i,\alpha}(\theta_{0})$ (Assumption 2.1) is fundamental for the proof of Asymptotic Normality because the correct specification ensures that the second term of the Hessian, $\mathbb{E}[\nabla g_{i}(\theta_{0})^{\top}r_{i}(\theta_{0})]$, vanishes because $\mathbb{E}[r_{i}(\theta_{0})\mid x_{i}]=0$. Without this, the ”bread” of the sandwich covariance matrix would require an additional complex term involving the second derivatives of the regression function, and the standard information matrix equality would fail.
2.4 Robust extensions of the $\alpha$–regression